OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (anonymous):

lol please help me in this question ;) lol

OpenStudy (anonymous):

i need help to answer this question so i know how to respond to the rest

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

dont give me the answer give me the walk through

OpenStudy (anonymous):

Let's assume that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50

OpenStudy (anonymous):

I think it about 50 human and 24 Horses

OpenStudy (anonymous):

I get medal ur Bestie Dakota

OpenStudy (anonymous):

oh thanks @Country_Boy_Jon1997. and @live10000000

OpenStudy (anonymous):

Of course I'm goin 2 help my best friend

OpenStudy (anonymous):

Let's assume that HM = Human and HR = Horse HM + HR = 74 2HM + 4HR = 196 (2HM + 4HR) - (2 HM + 2HR) = 196 - 148 2HR = 48 HR = 24 HM + (24) = 74 HM = 74 - 24 HM = 50

OpenStudy (anonymous):

omg im starting to get it now

OpenStudy (anonymous):

its like a puzzle

OpenStudy (anonymous):

yeah puzzle is the key 2 Math

OpenStudy (anonymous):

No chat please Angeleye

OpenStudy (dan815):

haha what a joker

OpenStudy (anonymous):

please do not write thing not related 2 the question

OpenStudy (anonymous):

lol

OpenStudy (dan815):